How much 30% acid solution should be be added to 10% acid solution to make 100mL of a 12% acid solution?
Answer by addingup(3677)

The total is 100ml.
Let's call the amount at 0.30 x. How much of the solution can be at 0.10? Well, 100-x, since the total is 100.
0.30x+0.10(100-x) = 0.12(100)
0.30x+10-0.10x = 12
0.20x = 2
x = 10ml of the 30% solution. And:
100-10 = 90ml of the 10% solution
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
common sense quick check:
We want 12% and we have 30% and 10%. Is 12 closer to 30 or closer to 10? It's closer to 10, so it makes sense that the amount at 10% will be greater.